OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 98 when x = 7. A. g(x) = 14x2 B. g(x) = 2x2 C. g(x) = 7x2 D. g(x) = ½x2

OpenStudy (anonymous):

@JjJustin

OpenStudy (anonymous):

g(x) = Cx^2 98 = C7^2 C = 98/7^2 C = 2 g(x) = 2x^2